OpenStudy (anonymous):

(fxg) i cant figure out the answer f=Squareroot of x-2 g= x+6

OpenStudy (curry):

\[\sqrt{x-2}+6\] dont forgot medal

OpenStudy (anonymous):

(fog)(x)=sqrt((x+6)-2) =sqrt(x+4)